I don't understand this problem. Could you please help!
Find the equation of each line and express in the form ax+by=c
Problem:
Line passes through (-5, -1) and is perpendicular to -1/3X-Y=2
Thanks SO MUCH!!

Answer by longjonsilver(2297)

line (-1/3)x-y=2 --> y = (-1/3)x-2, so its gradient is (-1/3). Any line perpendicular to this has gradient of 3 (2 perpendicular gradients, m*n = -1)

So we have x=-5, y=-1, m=3...find c from y=mx+c and that will give you the equation of the line.

jon.